On January 1, 2018, Ameen Company purchased major pieces of manufacturing equipment for a total of $68 million. Ameen uses straight-line depreciation for financial statement reporting and MACRS for income tax reporting. At December 31, 2020, the book value of the equipment was $62 million and its tax basis was $52 million. At December 31, 2021, the book value of the equipment was $60 million and its tax basis was $45 million. There were no other temporary differences and no permanent differences. Pretax accounting income for 2021 was $45 million.

Required:
1. Prepare the appropriate journal entry to record Ameen’s 2021 income taxes. Assume an income tax rate of 25%.
2. What is Ameen’s 2021 net income?
Answer #1

Ans:

1  

Debit

Credit

Calculation

Income tax expense

$ 11.25

=45*25%

       Deferred tax liability

$ 1.25

=(52-45)-(62-60)*25%

       Income tax payable

$ 10

2

Ameen’s 2021 net income = 45-11.25 = $33.75 million